OpenStudy (anonymous):

(2/5)2 .25= order of operation and that 2 is little beside the other 2

OpenStudy (anonymous):

(2/5)2 *.25 or (2/5)*2.25?

OpenStudy (anonymous):

first one..but its a little 2 outside the )

OpenStudy (anonymous):

\[\large(\frac{2}{5})^2 * 0.25?\]

OpenStudy (anonymous):

yes how do u type that to makem it a little 2

OpenStudy (anonymous):

In the equation box, you can just type(2/5)^2. That little 2 is called an exponent.

OpenStudy (anonymous):

and of course answer

OpenStudy (anonymous):

An exponent of 2 means multiply (2/5) by (2/5)

OpenStudy (anonymous):

ok ty now i know

OpenStudy (anonymous):

.25=1/4 \[(\frac{2}{5})^2*\frac{1}{4}=\frac{4}{25}*\frac{1}{4}=\frac{4}{100}=\frac{1}{25}=0.04\]

OpenStudy (anonymous):

ty your a life savor'